OpenStudy (anonymous):

Alice put $68,000 in two different accounts. One account pays 6.5% and the other 7% interest. If the total amount earned in one year is $4,535, how much did she invest in each account?

OpenStudy (anonymous):

Just solve the system of equations in 2 variables. .065x + .07y = 4535 x+y=68000

OpenStudy (anonymous):

How do I solve the first equation?

hero (hero):

It's a system of equations. You solve them together.

OpenStudy (anonymous):

You solve them both at once. By substitution

OpenStudy (anonymous):

.065x^2+.07y^2=72535???????????

OpenStudy (anonymous):

No

OpenStudy (anonymous):

did I do that correct?

OpenStudy (anonymous):

Am I way off?

OpenStudy (anonymous):

4535 * 7 4535 * 65???

OpenStudy (anonymous):

68000+4535=72535?

OpenStudy (agent0smith):

.065x + .07y = 4535 x+y=68000 use the second equation, x+y=68000, to rearrange for y=, to get: y = 68000-x Now, replace y in the first equation with that: \[\Large .065x + .07(68000 - x) = 4535\]

OpenStudy (anonymous):

.065x+476-.07x=4535??

OpenStudy (agent0smith):

if you meant... \[\Large .065x+4760-.07x=4535\]Then it looks good so far (you had a 476 not 4760). Now try solving that for x. It might help to first rearrange the order like so: \[\Large .065x-.07x +4760 =4535\]

OpenStudy (anonymous):

.06x-.07x=-225

OpenStudy (anonymous):

how does that look?

OpenStudy (agent0smith):

Well.. close, but now you dropped another number... \[\Large .065x-.07x=-225 \] Now just combine the x terms and find x.

OpenStudy (anonymous):

add .07 to each side? add .065 to each side?

OpenStudy (agent0smith):

No, combine like terms. For example, 2x + 3x = 5x, or 4x - 7x = -3x. Combine the 0.065x-0.07x to get...?

OpenStudy (anonymous):

.135

OpenStudy (agent0smith):

One is being subtracted from the other...

OpenStudy (anonymous):

.005=-225

OpenStudy (anonymous):

.005x=-225

OpenStudy (agent0smith):

0.065x - 0.07x doesn't equal 0.005x, but you're close... be careful with negatives.

OpenStudy (anonymous):

-.005x=-225

OpenStudy (agent0smith):

Correct. Now find x.

OpenStudy (anonymous):

=45000

OpenStudy (anonymous):

x=45000

OpenStudy (anonymous):

is 45000 6.5%??

OpenStudy (agent0smith):

.065x + .07y = 4535 yep, since it has the 0.065 in front.

OpenStudy (anonymous):

23000 is 7%??

OpenStudy (agent0smith):

Yep.